### v1

# Two-Point Stabilizer Rényi Entropy: a Computable Magic Proxy of Interacting Fermions

Abstract
Quantifying non-stabilizerness (“magic”) in interacting fermionic systems remains a formidable challenge, particularly for extracting high order correlations from quantum Monte Carlo simulations. In this Letter, we establish the two-point stabilizer Rényi entropy (SRE) and its mutual counterpart as robust, computationally accessible probes for detecting magic in diverse fermionic phases. By deriving local estimators suitable for advanced numerical methods, we demonstrate that these metrics effectively characterize quantum phase transitions: in the one-dimensional spinless $t$-$V$ model, they sharply identify the Luttinger liquid to charge density wave transition, while in the two-dimensional honeycomb lattice via determinant quantum Monte Carlo, they faithfully capture the critical exponents of the Gross-Neveu-Ising universality class. Furthermore, extending our analysis to the fractional quantum Hall regime, we unveil a non-trivial spatial texture of magic in the Laughlin state, revealing signatures of short-range exclusion correlations. Our results validate the two-point SRE as a versatile and sensitive diagnostic, forging a novel link between quantum resource theory, critical phenomena, and topological order in strongly correlated matter.
Introduction.— Quantum entanglement, a pivotal resource in quantum computation, has become central to assessing the complexity of simulating quantum states on classical hardware. It has been established that states with low entanglement are efficiently representable by matrix product states [1], underpinning the efficacy of the density matrix renormalization group (DMRG) in one dimension [2, 3, 4, 5, 6, 7, 8, 9]. In this context, entanglement is widely regarded as a prerequisite for achieving universal quantum computation [10]. Simultaneously, it serves as a powerful diagnostic in many-body physics [3, 11, 7, 12], characterizing exotic condensed matter phases, such as topological orders [13, 14, 15, 16, 17], as well as critical phenomena observable in both quantum [18, 8] and finite-temperature phase transitions [19, 20]. Nevertheless, entanglement alone is insufficient to fully capture the computational complexity of quantum states. For instance, highly entangled states generated by Clifford circuits—such as the Bell state—remain classically simulable.
Within the Clifford framework, the Gottesman-Knill theorem [21, 22] dictates the efficient classical simulation of stabilizer states, thereby attributing the origin of quantum computational advantage to non-Clifford operations, such as T-gates. Consequently, non-Clifford gates and the resulting magic states constitute the essential ingredients for universal quantum computation [23, 24, 25, 26, 27, 28]; this specific resource is formally termed quantum non-stabilizerness, or quantum magic. Recent developments highlight a non-trivial interplay between entanglement and magic in decoding the information structure of quantum states [29, 30, 31]. This synergy is particularly salient in the context of measurement-induced phase transitions [32], where distinct dynamical phases are classified by their respective scaling laws with system size.
To quantify such resources, a variety of measures have been introduced, including relative entropy of magic, mana [33, 34], and robustness of magic [35, 36, 37, 38]. However, evaluating these metrics becomes prohibitively expensive for large-scale systems. Addressing this limitation, the stabilizer Rényi entropy (SRE) has gained prominence as a computationally tractable and discriminative quantifier of magic [39, 40].
To efficiently evaluate the SRE, specialized numerical strategies have been devised to optimize the Pauli string summation in spin systems [41, 42, 43, 44, 45, 46, 47]. Conversely, the understanding of SRE in static fermionic contexts remains sparse, primarily restricted to free fermions [48, 49] and the SYK model [50, 51]. A general computational framework for large scale interacting fermions is conspicuously absent, especially in two dimensions where DMRG is severely impeded by entanglement scaling [52, 53].
In this Letter, we bridge this gap by establishing a determinant quantum Monte Carlo (DQMC) framework [54, 55, 56] to quantify non-stabilizerness in interacting fermionic systems. Crucially, to circumvent the excessive sampling overhead associated with global rank-$\alpha$ SRE, a bottleneck we verify against 1D DMRG benchmarks, we adopt the two-point SRE [57, 46, 58] as a robust and tractable proxy, complemented by the mutual two-point SRE to isolate correlation induced contributions. By deriving local estimators tailored for DQMC, we surmount the signal-to-noise limitations inherent to global measures. We apply this methodology to the half-filled spinless $t$-$V$ model on both one-dimensional chains and two-dimensional honeycomb lattices. Our investigation uncovers that the two-point SRE serves as a sharp detector of quantum phase transitions. Notably, via finite-size scaling analysis, we demonstrate that this magic based metric faithfully captures the critical exponents of the Gross-Neveu-Ising universality class, forging a novel connection between quantum magic and fermionic criticality. Finally, we extend our scope to fractional quantum Hall systems, specifically the Laughlin state, elucidating the spatial texture of non-stabilizerness by mapping magic correlations across orbital distances.
SRE for fermions in one dimension.— We begin by defining the rank-$\alpha$ SRE, $M_{\alpha}$, as
$M_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\sum_{P\in\mathcal{P}_{N}}\frac{1}{d}\left[\text{Tr}(P\rho)\right]^{\alpha}-S_{\alpha}(\rho),$ (1)
where $P$ runs over the $N$-qubit Pauli group $\mathcal{P}_{N}$, $d=2^{N}$ denotes the Hilbert space dimension, and $S_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\text{Tr}(\rho^{\alpha})$ is the rank-$\alpha$ Rényi entropy. In fermionic settings, Pauli strings are replaced by Majorana strings $\gamma^{\bm{x}}$, whose expectation values are efficiently evaluated for fermionic Gaussian states, allowing the sum over Majorana strings to be performed via Monte Carlo importance sampling [48, 49]. Since the density matrix of interacting fermions is approximated as a weighted ensemble of Gaussian states via Trotter decomposition and Hubbard-Stratonovich transformation, the evaluation of SRE becomes, in principle, feasible within the DQMC framework. We employ a two-level Monte Carlo sampling scheme by first sampling the auxiliary fields and subsequently the Majorana strings (see Supplemental Material (SM) [59] for details). In Fig. 1, we present the rank-2 SRE computed via DQMC for the one-dimensional half-filled $t$-$V$ model
$H=-t\sum_{\langle i,j\rangle}\left(c_{i}^{\dagger}c_{j}+\text{H. c.}\right)+V\sum_{\langle i,j\rangle}\left(n_{i}-\frac{1}{2}\right)\left(n_{j}-\frac{1}{2}\right)$ (2)
subject to periodic boundary conditions, benchmarking our results against DMRG calculations. Both methods indicate that the SRE exhibits volume law scaling [49, 60, 61], $M_{\alpha}\sim L$; however, the two-level Monte Carlo sampling scheme is computationally expensive, limiting access to large system sizes. This challenge arises from the high-order correlation functions inherent to the global SRE, which necessitate a large number of samples in the simulations.
Figure 1: Rank-2 SRE as a function of chain length $L$ for the one-dimensional half-filled $t$-$V$ model at various interaction strengths $V$ ($t=1$). Open and solid circles denote DQMC and DMRG results, respectively, while solid lines represent linear fits to the DMRG data. Both approaches confirm volume law scaling, $M_{2}\sim L$. The DQMC data for $L\geq 26$ exhibit large error bars due to insufficient sampling, as the high computational cost of the two-level Monte Carlo scheme limits access to large system sizes.
Figure 2: Scaling behavior of the mutual two-point SRE in the one-dimensional half-filled spinless $t$-$V$ model. (a) Distance dependence of the mutual SRE in the LL phase (blue circles) and CDW phase (red squares) for $L=400$. Interaction strengths are chosen deep within each phase to minimize finite size effects, clearly exhibiting algebraic and exponential decay, respectively. (b) Mutual two-point SRE as a function of interaction strength $V$ for $L=900$ at various distances. A pronounced peak identifies the critical transition. (c) Finite size scaling of the critical points $V_{c}(L)$, it obeys the scaling relation predicted by BKT transition. The non-universal microscopic length scale $L_{0}$ is fitted to be $L_{0}\approx 1.7$ with data $L>100$ are considered in the fitting.
Two-point SRE in one dimension.— To bypass the aforementioned sampling hurdles, we adopt the two-point SRE [57, 46, 58], defined as $\mathcal{M}^{(\alpha)}_{i,j}(\rho)=M_{\alpha}(\rho_{i,j})$, where $\rho_{i,j}$ denotes the two site reduced density matrix. Since the partial trace operation is monotonically non-increasing regarding non-stabilizerness [39], this quantity establishes a lower bound for the global magic, i.e., $\mathcal{M}^{(\alpha)}_{i,j}(\rho)\leq M_{\alpha}(\rho)$. While $\mathcal{M}^{(\alpha)}_{i,j}$ vanishes for all stabilizer states, the converse does not strictly hold, rendering it a globally non-faithful measure. However, a non-zero value definitively signals the presence of non-Clifford resources, thereby classifying it as a valid magic witness. Despite the lack of global faithfulness, it constitutes a potent probe for local “magic density” and correlations. Given the local nature of the Hamiltonian in our study (e.g., the $t$-$V$ model), non-stabilizerness is anticipated to manifest locally. Consequently, the two-point SRE functions as an “effective order parameter”, detecting phase transitions characterized by the emergence of local non-Clifford resources and linking abstract quantum information metrics with accessible experimental observables.
We revisit the one-dimensional half-filled $t$-$V$ model [62, 63, 64], utilizing DMRG simulations [65] with open boundary conditions (centering the reference site $i$). We first analyze the spatial scaling behavior as the distance |ij||i-j| varies. According to bosonization theory, the single particle Green’s function $\langle c_{i}^{\dagger}c_{i+r}\rangle$ decays algebraically as $r^{-\eta}$ (with $\eta\approx 1$) in the Luttinger liquid (LL) phase ($V<2$), whereas it exhibits exponential decay $e^{-r/\xi}$ in the gapped charge density wave (CDW) phase. Leveraging tomography techniques [66], we derive the general analytical form of the fermionic two-point SRE. To a first order approximation, this quantity is dominated by the squared magnitude of the Green’s function, superimposed onto a background of intrinsic magic density. In the LL regime, the single site reduced density matrix corresponds to a mixed stabilizer state, yielding zero intrinsic magic. In contrast, the CDW phase retains a finite magic density arising from spontaneous symmetry breaking. To isolate the correlation dependent scaling from this local background, we introduce the mutual two-point SRE, defined as $\tilde{\mathcal{M}}^{(\alpha)}_{i,j}(\rho)=\mathcal{M}_{i,j}^{(\alpha)}(\rho)-\mathcal{M}_{i}^{(\alpha)}(\rho)-\mathcal{M}_{j}^{(\alpha)}(\rho)$. Figure 2(a) displays the computed mutual two-point SRE across distinct phases. The numerical results corroborate our analytical predictions. Specifically, in the interacting LL phase, we observe that the power law decay exponent deviates slightly from the non-interacting limit of $2$, reflecting interaction effects. Subsequently, we confirm that this exponent modification, and the correlation length $\xi(V)$, can be precisely extracted via the susceptibility $\partial\tilde{\mathcal{M}}/\partial V$, providing an effective scheme for characterizing the phase transition.
Upon tuning the interaction strength $V$, $\tilde{\mathcal{M}}^{(2)}_{i,j}$ emerges as a sensitive indicator of the quantum phase transition between the LL and CDW phases. As illustrated in Fig. 2(b), the mutual two-point SRE displays a pronounced peak characterized by a discontinuity in its first derivative at the critical point, a feature that becomes sharper as the system size increases. We attribute this behavior to the distinct scaling laws governing $\tilde{\mathcal{M}}$ in the respective phases. In the gapless LL regime, the algebraic decay yields a derivative scaling of $\frac{\partial\mathcal{M}_{i,i+r}^{(2)}}{\partial V}\sim-2\frac{\partial\alpha}{\partial V}r^{-2\alpha(V)}\ln r$. Conversely, in the gapped CDW phase, the exponential decay leads to $\frac{\partial\mathcal{M}_{i,i+r}^{(2)}}{\partial V}\sim\frac{r}{\xi^{2}(V)}e^{-r/\xi(V)}\frac{\partial\xi}{\partial V}$. We identify the location of this maximum as the finite-size critical point, $V_{c}(L)$. Notably, the shift of $V_{c}(L)$ from the asymptotic value $V_{c}(\infty)=2$ follows the characteristic finite size scaling of a Berezinskii-Kosterlitz-Thouless (BKT) transition [67, 68, 69, 70, 71, 72], where the correlation length diverges as $\xi\sim\exp{\left(\pi^{2}/2\sqrt{|V-V_{c}|}\right)}$. Equating the correlation length to the system size, $\xi\sim L$, yields the scaling relation $V_{c}(L)=2+\frac{\pi^{4}/4}{(\ln L-\ln L_{0})^{2}}$ where $L_{0}$ is a non-universal microscopic length scale. As shown in Fig. 2(c), $V_{c}(L)$ is in excellent agreement with this scaling relation. Therefore, we conclude that the two-point SRE accurately captures the critical scaling properties of the transition.
Two-point SRE in two dimensions.— We now extend our analysis to the honeycomb lattice implementation of Eq. (2). This system exhibits a quantum phase transition from a Dirac semimetal (DSM) to a CDW phase at $V_{c}\approx 1.355$ [73, 74]. Here, we employ DQMC to compute the mutual two-point SRE. As shown in Fig. 3(a), near criticality, $\tilde{\mathcal{M}}^{(2)}_{i,i+\vec{r}_{\max}}$ (with $\vec{r}_{\max}=(L/2,L/2)$) exhibits a power law decay with increasing system size $L$. We attribute this behavior to the fact that, in the vicinity of the critical point, $\tilde{\mathcal{M}}^{(2)}$ is governed by the squared density-density correlation $C_{\mathrm{max}}(L)=\langle(n_{i}-1/2)(n_{i+\vec{r}_{\mathrm{max}}}-1/2)\rangle$. Given that this correlator scales as $L^{-1-\eta}$ [74], where $\eta$ represents the bosonic anomalous dimension, the exponent can be extracted directly from the SRE scaling. According to finite-size scaling theory, the density correlation follows the ansatz [75]
$C_{\mathrm{max}}(L)=L^{-1-\eta}\mathcal{F}(L^{1/\nu}(V-V_{c})).$ (3)
To circumvent the uncertainty in $V_{c}$, we perform the scaling analysis against the dimensionless correlation ratio $R=1-\frac{C(k=q_{\text{min}})}{C(k=0)}$ [76], where $C(k)$ is the Fourier transform of the density-density correlation and $q_{\text{min}}=(2\pi/L,0)$. Using the package in Ref. [77], we determine $\eta\approx 0.423$, which is consistent with Gross-Neveu-Ising theory [78], achieving optimal data collapse for the scaled $\tilde{\mathcal{M}}^{(2)}_{i,i+\vec{r}_{\max}}$ versus $R$, as shown in Fig. 3(b). This result further confirms the capability of the mutual two-point SRE to accurately determine critical exponents at two-dimensional quantum critical points.
Orbital texture of magic in FQH states.— We investigate the orbital architecture of the mutual SRE in the $\nu=1/3$ Laughlin state [79]. Imposed by rotational symmetry, the two-orbital reduced density matrix is strictly diagonal, and thus uniquely determined by the joint occupation probability $P_{11}$. As illustrated in Fig. 4, at short range, we observe a distinct plateau where the mutual magic remains constant. In the language of Haldane pseudopotentials, this signifies that the state possesses a vanishing projection onto relative angular momentum channels with $L_{\text{rel}}<m$ (specifically excluding $L_{\text{rel}}=1$ and $2$ for fermions at $\nu=1/3$) [80, 81]. Consequently, fixing a particle at orbital 0, the occupation probability for intermediate orbitals $j<m$ is rigorously suppressed. Within the resource theory framework, this angular momentum constraint locks the local state into a regime of maximal non-stabilizerness allowed by the filling factor.
Figure 3: Critical scaling on the honeycomb lattice. (a) Mutual two-point SRE $\tilde{\mathcal{M}}^{(2)}_{i,i+\vec{r}_{\max}}$ as a function of linear system size $L$ at various interaction strengths $V$ near the critical threshold ($V_{c}\approx 1.34$ [82]). The dashed line indicates a power-law fit characteristic of the critical point. (b) Finite-size scaling collapse of the rescaled SRE, $\tilde{\mathcal{M}}^{(2)}L^{2+2\eta}$, plotted against the correlation ratio $R$. The optimal data collapse yields an anomalous dimension $\eta\approx 0.423$. Note that data for $V>1.36$ are excluded from the fit to ensure a robust estimation of $\eta$.
Following this exclusion-induced plateau, the mutual SRE exhibits a precipitous decline until the joint occupation $P_{11}$ surpasses the “magic critical threshold” $x_{c}=\nu_{\text{eff}}-1/4$. Above this threshold, the mutual SRE synchronously tracks $P_{11}$, mirroring its spatial oscillations before asymptotically decaying to zero at large separations [83]. This behavior encapsulates the characteristic exponential decay inherent to a gapped quantum liquid. Collectively, these results establish the two-point SRE as a potent diagnostic for resolving the non-trivial internal texture of fractional quantum Hall states.
Conclusion and outlook.— We have established that the global SRE in fermionic systems is generally computationally intractable due to the complexity of high order correlation functions. To surmount this hurdle, we championed the two-point SRE as a powerful local probe for identifying “magic density” and correlations. By introducing the mutual two-point SRE, we successfully isolated intrinsic magic from background contributions. Utilizing DMRG for the one-dimensional half-filled spinless $t$-$V$ model, we demonstrated that this metric acts as a sharp detector of the Luttinger liquid to CDW transition. Furthermore, employing our DQMC framework on the honeycomb lattice, we verified that the mutual SRE accurately resolves the critical exponents of the Gross-Neveu-Ising universality class. Expanding beyond lattice models, our analysis of the Laughlin state elucidated the spatial texture of non-stabilizerness in topological fluids, revealing how the Pauli exclusion principle dictates the short-range distribution of magic. Collectively, these findings forge a novel link between quantum magic and fermionic criticality, validating the two-point SRE as a versatile diagnostic for strongly correlated matter.
Figure 4: Orbital texture of non-stabilizerness in the $\nu=1/3$ Laughlin liquid ($N=9$). The mutual two-point SRE (red circles) is plotted against the orbital distance $j$, alongside the two-body density correlation $P_{11}=\langle n_{0}n_{j}\rangle$ (blue squares). The green dot-dashed line marks the magic critical threshold $x_{c}=\nu_{\text{eff}}-1/4$. Note the pronounced plateau at short range ($j=1,2$) arising from the vanishing component of relative angular momentum $L_{\text{rel}}<m$, and the subsequent synchronization between magic revival and density modulation at intermediate distances.
While this study focused on ground state properties, the developed DQMC protocols are inherently applicable to finite temperature Gibbs states. Consequently, exploring the thermal evolution of rank-$\alpha$ SRE across phase transitions represents a natural progression. Future avenues also include generalizations to multi point SRE and spinful fermionic models—incorporating spin-orbit couplings—which promise to unveil richer tapestries of symmetry-resolved magic and intricate non-Clifford correlations in many-body systems.
Acknowledgments.— We thank Chenghao He for fruitful discussion. This work was supported by the National Natural Science Foundation of China (Grants No. 12447103, No. 12274289 and No. 125B2077), the National Key R&D Program of China (Grants No. 2022YFA1402702 and No. 2021YFA1401400), the Innovation Program for Quantum Science and Technology (under Grant No. 2021ZD0301902), the Yangyang Development Fund, and Shanghai Jiao Tong University 2030 Initiative. F.-H. W. is supported by T.D. Lee scholarship. The computations in this paper were run on the Siyuan-1 and $\pi$ 2.0 clusters supported by the Center for High Performance Computing at Shanghai Jiao Tong University.
Supplemental Material for “Two-point stabilizer Rényi entropy: a computable magic proxy of interacting fermions via quantum Monte Carlo simulations”
Jun Qi Fang1, Fo-Hong Wang1,2 and Xiao Yan Xu1,3,∗
1Key Laboratory of Artificial Structures and Quantum Control (Ministry of Education),
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
2Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 200240, China
3Hefei National Laboratory, Hefei 230088, China
In this Supplementary Material, we provide the theoretical foundations and technical details supporting the findings presented in the main text. In Sec. SI, we derive the formalism for evaluating the Stabilizer Rényi Entropy (SRE) in fermionic Gaussian states via the Majorana monomial mapping and outline the implementation of our determinant quantum Monte Carlo (DQMC) algorithm for interacting systems. In Sec. SII, we detail the quantum state tomography protocol employed to reconstruct two-point reduced density matrices from local observables. Furthermore, we present comprehensive scaling analyses for the spinless $t$-$V$ model in both one and two dimensions, as well as for the Laughlin state, substantiating the phase transition criteria and critical exponents discussed in the main text.
SI SRE in fermionic Gaussian states and DQMC simulation
In this section, we detail the calculation of the SRE for fermionic Gaussian states and outline the DQMC scheme employed for interacting fermionic systems.
We begin by illustrating the procedure for determining the SRE of a fermionic Gaussian state, which serves as a crucial component of our DQMC framework. As established in Refs. [48, 49], the summation over Pauli strings can be mapped to a summation over Majorana monomials. This transformation enables the efficient evaluation of SRE for fermionic Gaussian states. Specifically, the sum over the Pauli group is equivalent to the sum over Majorana monomials:
$\sum_{P\in\mathcal{P}_{n}}\text{Tr}[P\rho]^{2\alpha}=\sum_{\gamma^{\bm{x}}}\left|\text{Tr}[\gamma^{\bm{x}}\rho]\right|^{2\alpha},$ (S1)
where $\hat{\gamma}^{\bm{x}}=(\hat{\gamma}_{1})^{x_{1}}(\hat{\gamma}_{2})^{x_{2}}\cdots(\hat{\gamma}_{2N})^{x_{2N}}$ denotes a Majorana monomial, with the summation running over all binary configurations $\bm{x}\in\{0,1\}^{2N}$, and $N$ representing the number of sites. These Majorana operators are constructed via the standard Jordan-Wigner transformation:
$\hat{\gamma}_{2i-1}$ $=\hat{Z}_{1}.\hat{Z}_{i-1}\hat{X}_{i}\hat{1}_{i+1}.\hat{1}_{N}$ (S2)
$\hat{\gamma}_{2i}$ $=\hat{Z}_{1}.\hat{Z}_{i-1}\hat{Y}_{i}\hat{1}_{i+1}.\hat{1}_{N}.$
They satisfy the anti-commutation relations $\{\hat{\gamma}_{\mu},\hat{\gamma}_{\nu}\}=2\delta_{\mu\nu}$. Here, the Pauli operators involved are $\hat{\mathds{1}},\hat{X}_{i},\hat{Y}_{i}$, and $\hat{Z}_{i}$. Equivalently, these operators can be expressed in terms of complex fermionic operators $c_{i}$ and $c_{i}^{\dagger}$ as $\hat{\gamma}_{2i-1}=(\hat{c}_{i}^{\dagger}+\hat{c}_{i})$ and $\hat{\gamma}_{2i}=i(\hat{c}_{i}^{\dagger}-\hat{c}_{i})$.
The validity of Eq. (S1) rests on the bijective relationship between Majorana monomials and Pauli strings. Let $\hat{\tilde{P}}_{i}=\hat{\mathds{1}}_{1}\cdots\hat{\mathds{1}}_{i-1}\hat{P}_{i}\hat{\mathds{1}}_{i+1}\cdots\hat{\mathds{1}}_{N}$, where $\hat{P}_{i}\in\{\hat{\mathds{1}}_{i},\hat{X}_{i},\hat{Y}_{i},\hat{Z}_{i}\}$ acts on a single site. A generic Pauli string is given by $\bm{P}=\prod_{i=1}^{N}\hat{\tilde{P}}_{i}$. Using the identities $\hat{\gamma}_{2i-1}\hat{\gamma}_{2i}=i\hat{\tilde{Z}}_{i}$ and $\hat{\gamma}_{\mu}^{2}=\hat{\mathds{1}}$, the operators $\hat{\tilde{X}}_{i}$ and $\hat{\tilde{Y}}_{i}$ can be reconstructed as:
$\tilde{X}_{i}=\left(\prod_{j=1}^{i-1}\tilde{Z}_{j}\right)\gamma_{2i-1}=\left[\prod_{j=1}^{i-1}(-i)\gamma_{2j-1}\gamma_{2j}\right]\gamma_{2i-1}$ (S3)
$\tilde{Y}_{i}=\left(\prod_{j=1}^{i-1}\tilde{Z}_{j}\right)\gamma_{2i}=\left[\prod_{j=1}^{i-1}(-i)\gamma_{2j-1}\gamma_{2j}\right]\gamma_{2i}.$
Consequently, the relationship between an arbitrary Pauli string $\hat{\bm{P}}$ and a Majorana monomial $\hat{\gamma}^{\bm{x}}$ is given by $\hat{\bm{P}}=e^{i\phi}\hat{\gamma}^{\bm{x}}$. Since the SRE depends on $\text{Tr}(\bm{P}\rho)]^{2}$, the phase factor $e^{i\phi}$ is irrelevant, making the summation over Majorana monomials equivalent to that over Pauli strings but computationally more straightforward.
For a Gaussian state $\rho_{G}$, Wick’s theorem allows the expectation value of $\hat{\gamma}^{\bm{x}}$ to be computed as a Pfaffian:
$\text{Tr}(\gamma^{\bm{x}}\rho_{G})=\text{Pf}[\Gamma|_{\bm{x}}],$ (S4)
where $\Gamma_{kl}=\frac{1}{2}\langle[\gamma_{k},\gamma_{l}]\rangle_{\rho_{G}}$ is the covariance matrix of $\rho_{G}$ (following the convention of Ref. [49]). Here, $\Gamma|_{\bm{x}}$ denotes the sub-matrix of $\Gamma$ formed by retaining only the rows and columns corresponding to indices where $x_{k}=1$. Accordingly, the SRE takes the form:
$M_{\alpha}(\rho_{G})=\frac{1}{1-\alpha}\log\sum_{\hat{\gamma}^{\bm{x}}}\pi_{\rho_{G}}^{\alpha}(\bm{x}),$ (S5)
where we defined $\pi_{\rho_{G}}(\bm{x})=\frac{\det[\Gamma|_{\bm{x}}]}{\det[1+\Gamma]}$. Note that for pure states, which are the focus here, $S_{\alpha}(\rho_{G})=0$. The calculation of rank-$\alpha$ SRE thus reduces to evaluating the mean of $\pi^{\alpha-1}_{\rho_{G}}(\bm{x})$ sampled from the probability distribution $\pi_{\rho_{G}}(\bm{x})$. To this end, we employ the Majorana sampling method introduced in Ref. [49] to efficiently generate the relevant configurations. We note that this scheme is also directly extendable to mixed Gaussian states.
We now turn to interacting fermionic systems, where the density matrix is inherently non-Gaussian. Within the DQMC framework, the Trotter-Suzuki decomposition and Hubbard-Stratonovich transformation decouple the interaction terms into fermion bilinears coupled to fluctuating auxiliary fields $\bm{s}$ in spacetime. The resulting partition function is $Z=\sum_{\mathbf{s}}\mathrm{Tr}[\prod_{l=1}^{L_{\tau}}e^{\mathbf{c}^{\dagger}K_{l}[\mathbf{s}]\mathbf{c}}]$, where $L_{\tau}$ denotes the number of imaginary-time slices and $\mathbf{c}=(c_{1},\cdots,c_{N})^{T}$ (spin indices omitted for brevity). Consequently, the density matrix $\rho$ can be expressed as a weighted ensemble of Gaussian operators, $\rho=\sum_{\bm{s}}P_{\bm{s}}\rho_{\bm{s}}$, where $P_{\bm{s}}$ represents the statistical weight of the auxiliary field configuration $\bm{s}$. Substituting this decomposition into Eq. (S1), the evaluation reduces to computing the numerator:
$\text{Tr}(\rho\gamma^{\bm{x}})=\sum_{\bm{s}}P_{\bm{s}}\text{Tr}(\rho_{\bm{s}}\gamma^{\bm{x}})=\sum_{\bm{s}}P_{\bm{s}}\text{Pf}(\Gamma_{\bm{s}}|_{\bm{x}}).$ (S6)
To efficiently sample Majorana monomials [50], we benchmarked various update strategies, identifying the Pauli string update as the most efficient. The procedure involves initializing a Pauli string $\bm{P}$ and proposing a new configuration $\bm{P}^{\prime}$ by randomly modifying operators at two distinct sites, which is subsequently mapped back to the Majorana representation. Due to the non-local nature of the Jordan-Wigner string, this operation constitutes a global update in the Majorana basis, significantly suppressing autocorrelation times.
SII Tomography technique for two-point SRE and scaling analysis
In this section, we detail the quantum state tomography technique employed to construct the two-point reduced density matrix, $\rho_{i_{1}i_{2}}$, and present the scaling analysis for both the $t$-$V$ model and the Laughlin state. Following the protocol established in Ref. [66], we derive the general form of $\rho_{i_{1}i_{2}}$ for systems with broken translational symmetry (e.g., open boundary conditions in DMRG). This reconstruction requires measuring a complete set of local observables, specifically the single-particle Green’s function and density-density correlations.
We assume the basis is ordered as $\{|0\rangle,c_{i_{2}}^{\dagger}|0\rangle,c_{i_{1}}^{\dagger}|0\rangle,c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle\}$ (where indices $i_{1}$ and $i_{2}$ correspond to the two sites in question, with $i_{1}$ typically being the reference site). The off-diagonal elements are determined as follows:
$[\rho_{i_{1}i_{2}}]_{23}$ $=\langle 0|c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}|0\rangle$ (S7)
$=\langle 0|c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}|0\rangle+\underbrace{\langle 0|c_{i_{2}}c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle}_{=0}+\underbrace{\langle 0|c_{i_{1}}c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{1}}^{\dagger}|0\rangle}_{=0}+\underbrace{\langle 0|c_{i_{2}}c_{i_{1}}c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle}_{=0}$
$=\mathrm{Tr}(\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}})$
$=\langle c_{i_{1}}^{\dagger}c_{i_{2}}\rangle\equiv g_{r},$
and similarly we have $\rho_{i_{1}i_{2}}]_{32}=g_{r}^{*}$. The diagonal elements are computed through analogous arguments involving density operators. For the joint occupation component $\rho_{i_{1}i_{2}}]_{44}$:
$[\rho_{i_{1}i_{2}}]_{44}$ $=\langle 0|c_{i_{2}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle$ (S8)
$=\langle 0|c_{i_{2}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle+\underbrace{\langle 0|c_{i_{2}}c_{i_{2}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle}_{=0}$
$\ \ \ +\underbrace{\langle 0|c_{i_{1}}c_{i_{2}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}c_{i_{1}}^{\dagger}|0\rangle}_{=0}+\underbrace{\langle 0|c_{i_{2}}c_{i_{1}}c_{i_{2}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle}_{=0}$
$=\text{Tr}(c_{i_{2}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger})$
$=\mathrm{Tr}(\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{1}}c_{i_{2}}^{\dagger}c_{i_{2}})$
$=\langle n_{i_{1}}n_{i_{2}}\rangle\equiv\varrho_{r}^{\mathrm{n}}.$
The single-particle occupancy elements are given by:
$[\rho_{i_{1}i_{2}}]_{33}$ $=\langle 0|c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}|0\rangle$ (S9)
$=\langle 0|c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}|0\rangle+\langle 0|c_{i_{2}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle+\underbrace{\langle 0|c_{i_{1}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{1}}^{\dagger}|0\rangle}_{=0}$
$\ \ \ +\underbrace{\langle 0|c_{i_{2}}c_{i_{1}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle}_{=0}-\langle 0|c_{i_{2}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle$
$=\mathrm{Tr}(\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{1}})-\langle 0|c_{i_{2}}c_{i_{1}}\rho_{i_{1}i_{2}}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle$
$=\langle n_{i_{1}}\rangle-\langle n_{i_{1}}n_{i_{2}}\rangle$
$[\rho_{i_{1}i_{2}}]_{22}$ $=\langle 0|c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{2}}^{\dagger}|0\rangle$ (S10)
$=\langle 0|c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{2}}^{\dagger}|0\rangle+\underbrace{\langle 0|c_{i_{2}}c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{2}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle}_{=0}+\langle 0|c_{i_{1}}c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{2}}^{\dagger}c_{i_{1}}^{\dagger}|0\rangle$
$\ \ \ +\underbrace{\langle 0|c_{i_{2}}c_{i_{1}}c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{2}}^{\dagger}c_{i_{1}}^{\dagger}c_{i_{2}}^{\dagger}|0\rangle}_{=0}-\langle 0|c_{i_{1}}c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{2}}^{\dagger}c_{i_{1}}^{\dagger}|0\rangle$
$=\mathrm{Tr}(\rho_{i_{1}i_{2}}c_{i_{2}}^{\dagger}c_{i_{2}})-\langle 0|c_{i_{1}}c_{i_{2}}\rho_{i_{1}i_{2}}c_{i_{2}}^{\dagger}c_{i_{1}}^{\dagger}|0\rangle$
$=\langle n_{i_{2}}\rangle-\langle n_{i_{1}}n_{i_{2}}\rangle$
Finally, the vacuum element $\rho_{i_{1}i_{2}}]_{11}$ is determined by the normalization condition $\text{Tr}(\rho_{i_{1}i_{2}})=1$, yielding $\rho_{i_{1}i_{2}}]_{11}=1-\langle n_{i_{1}}\rangle-\langle n_{i_{2}}\rangle+\langle n_{i_{1}}n_{i_{2}}\rangle$. All other matrix elements vanish.
SII.1 Scaling Analysis
We now proceed to the scaling analysis. By substituting the tomography relations Eqs. (S8)-(S10) into the definition of the rank-2 two point SRE, $\mathcal{M}^{(2)}_{i,j}=M_{2}(\rho_{i,j})$, the contribution arising from the Majorana string summation can be expanded in terms of local observables as
$\sum_{\gamma^{\bm{x}}}\frac{1}{d}\left|\text{Tr}(\rho_{i,j}\gamma^{\bm{x}})\right|^{4}=\frac{1}{4}(1+m_{i}^{4}+m_{j}^{4}+m_{i}^{4}m_{j}^{4}+32g_{I}^{4}+32g_{R}^{4}+256\delta^{4}+256m_{i}m_{j}\delta^{3}+96m_{i}^{2}m_{j}^{2}\delta^{2}+16m_{i}^{3}m_{j}^{3}\delta),$ (S11)
where $\gamma^{\bm{x}}=\gamma_{2i-1}^{n_{2i-1}}\gamma_{2i}^{n_{2i}}\gamma_{2j-1}^{n_{2j-1}}\gamma_{2j}^{n_{2j}}$ represents the Majorana string on the two sites, with the exponents $\{n_{2i-1},n_{2i},n_{2j-1},n_{2j}\}$ spanning the binary configurations $\{0,1\}^{4}$. The physical parameters are defined as the local magnetization $m_{i}=1-2\langle n_{i}\rangle$, the connected correlation $\delta=\langle n_{i}n_{j}\rangle-\langle n_{i}\rangle\langle n_{j}\rangle$, and the real and imaginary parts of the Green’s function, $g_{r}=g_{R}+ig_{I}$. Similarly, the purity term (entropy part) is:
$\text{Tr}(\rho_{i,j}^{2})=\frac{1+m_{i}^{2}}{2}\cdot\frac{1+m_{j}^{2}}{2}+2(g_{I}^{2}+g_{R}^{2})+2m_{i}m_{j}\cdot\delta+4\delta^{2}.$ (S12)
Assuming that correlations decay at long distances while local magnetizations $m_{i}$ remain $\mathcal{O}(1)$, a first order expansion of the two-point SRE yields:
$\ln\frac{(1+m_{i}^{2})(1+m_{j}^{2})}{(1+m_{i}^{4})(1+m_{j}^{4})}-\frac{32(g_{I}^{4}+g_{R}^{4})+16m_{i}^{3}m_{j}^{3}\delta+96m_{i}^{2}m_{j}^{2}\delta^{2}+256m_{i}m_{j}\delta^{3}+256\delta^{4}}{(1+m_{i}^{4})(1+m_{j}^{4})}+\frac{4(2(g_{I}^{2}+g_{R}^{2})+2m_{i}m_{j}\delta+4\delta^{2})}{(1+m_{i}^{2})(1+m_{j}^{2})}$ (S13)
Crucially, the constant background term $\ln\frac{(1+m_{i}^{2})(1+m_{j}^{2})}{(1+m_{i}^{4})(1+m_{j}^{4})}$ cancels out exactly in the mutual two-point SRE definition, leaving a quantity dominated by correlations.
1D $t$-$V$ Model: In the Luttinger liquid (LL) phase at half-filling ($n_{i}=0.5,m_{i}=0$), the mutual SRE is dominated by the Green’s function, scaling as $\tilde{\mathcal{M}}^{(2)}\sim|g_{r}|^{2}\sim r^{-2\eta}$ (where $|g_{r}|\sim r^{-\eta}$). In the CDW phase, the energy gap dictates exponential decay, $\tilde{\mathcal{M}}^{(2)}\sim e^{-r/\xi}$. The distinct scaling of the susceptibility $\partial_{V}\tilde{\mathcal{M}}^{(2)}$—diverging as $2\partial_{V}\alpha r^{-2\alpha}\ln r$ in the LL phase versus peaking as $\xi^{-2}re^{-r/\xi}\partial_{V}\xi$ in the CDW phase—signals the phase transition. Finite-size effects shift the peak of the latter to $V_{c}(L)$, governed by the BKT-like correlation length scaling $\xi\sim\exp(\pi^{2}/2\sqrt{|V-V_{c}|})$, leading to $V_{c}(L)\approx 2+\pi^{4}/4(\ln L-\ln L_{0})^{2}$.
2D Honeycomb $t$-$V$ Model: Near the critical point, symmetry ensures $m_{i}\approx 0$. Unlike the 1D case, here the mutual SRE is significantly influenced by density correlations $\delta$. Finite size scaling theory predicts that the connected density correlation scales as $\delta\sim L^{-(d+z-2+\eta)}$, while the Green’s function scales as $g_{r}\sim L^{-(2+\eta_{\psi})}$. In our case ($d=2,z=1$), the density term dominates at long distances, implying $\tilde{\mathcal{M}}^{(2)}\sim\delta^{2}\sim r^{-2(1+\eta)}$. This allows for the direct extraction of the anomalous dimension $\eta$.
Laughlin State: The $\nu=1/m$ Laughlin wavefunction for $N$ particles is given by [79]
$\Psi_{m}(z_{1},z_{2},...,z_{N})=\prod_{j<k}(z_{j}-z_{k})^{m}\exp\left(-\frac{1}{4}\sum_{i=1}^{N}|z_{i}|^{2}\right).$ (S14)
Defining the single-particle orbitals in the lowest Landau level as $\phi_{k}(z)\sim z^{k}e^{-|z|^{2}/4}$, which corresponds to the creation operator $c_{k}^{\dagger}\left|0\right\rangle$ in the Fock space [81], a generic many-body state is represented by a Slater determinant $\det[\phi_{k_{i}}(z_{j})]$. Consequently, the Laughlin wavefunction can be expanded in this occupation-number basis (or orbital basis) as
$\Psi_{m}(z_{1},z_{2},...,z_{N})=\sum_{\mathrm{config~}\{k_{1},\ldots,k_{N}\}}C^{m}_{\{k_{i}\}}\cdot\det[\phi_{k_{i}}(z_{j})].$ (S15)
Imposed by angular momentum conservation, the two-point reduced density matrix is strictly diagonal in this basis, $\rho_{0j}=\text{diag}(P_{00},P_{01},P_{10},P_{11})$, as the off-diagonal Green’s function $g_{r}$ vanishes. Here, the joint occupation probability relates directly to the radial distribution function via $P_{11}=\nu_{\text{eff}}^{2}g(r_{0j})$. The remaining elements follow from the tomography relations [Eqs. (S8)-(S10)]: $P_{10}=P_{01}=\nu_{\text{eff}}(1-\nu_{\text{eff}}g(r_{0j}))$, with $P_{00}$ fixed by normalization. We account for finite-size effects by employing the effective filling factor $\nu_{\text{eff}}=N/[m(N-1)+1]$. At short distances ($j<m$), the simultaneous occupation of the reference orbital 0 and target orbital $j$ is strictly forbidden, implying $g(r_{0j})=0$ (and thus $P_{11}=0$). This results in a constant mutual SRE plateau determined solely by $\nu_{\text{eff}}$. We further corroborate this plateau structure using the $\nu=1/5$ state, as shown in Fig. S1. Defining the variable $x\equiv P_{11}$, we identify a critical threshold $x_{c}=\nu_{\text{eff}}-1/4$. The mutual SRE decreases with $x$ in the regime $x<x_{c}$ and increases for $x>x_{c}$. Consequently, beyond the correlation hole, the oscillations of $\tilde{\mathcal{M}}^{(2)}$ synchronously track those of $P_{11}$, before asymptotically decaying to zero.
Figure S1: Analogous to Fig. 4 for the $\nu=1/5$ Laughlin state with $N=6$ particles. The mutual two-point SRE displays an extended plateau at short ranges ($j<5$). At intermediate distances, the magic signal synchronously tracks the oscillations of the joint occupation probability $P_{11}$.

### v2

@@
 SRE for fermions in one dimension.— We begin by defining the rank-$\alpha$ SRE, $M_{\alpha}$, as
-$M_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\sum_{P\in\mathcal{P}_{N}}\frac{1}{d}\left[\text{Tr}(P\rho)\right]^{\alpha}-S_{\alpha}(\rho),$ (1)
+$M_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\sum_{P\in\mathcal{P}_{N}}\frac{1}{d}\left[\text{Tr}(P\rho)\right]^{2\alpha}-S_{\alpha}(\rho),$ (1)
 where $P$ runs over the $N$-qubit Pauli group $\mathcal{P}_{N}$, $d=2^{N}$ denotes the Hilbert space dimension, and $S_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\text{Tr}(\rho^{\alpha})$ is the rank-$\alpha$ Rényi entropy.
@@
 Specifically, in the interacting LL phase, we observe that the power law decay exponent deviates slightly from the non-interacting limit of $2$, reflecting interaction effects.
-Subsequently, we confirm that this exponent modification, and the correlation length $\xi(V)$, can be precisely extracted via the susceptibility $\partial\tilde{\mathcal{M}}/\partial V$, providing an effective scheme for characterizing the phase transition.
+Subsequently, we confirm that this exponent modification, and the correlation length $\xi$, can be precisely extracted via the susceptibility $\partial\tilde{\mathcal{M}}/\partial V$, providing an effective scheme for characterizing the phase transition.
 Upon tuning the interaction strength $V$, $\tilde{\mathcal{M}}^{(2)}_{i,j}$ emerges as a sensitive indicator of the quantum phase transition between the LL and CDW phases.
@@
 We attribute this behavior to the distinct scaling laws governing $\tilde{\mathcal{M}}$ in the respective phases.
-In the gapless LL regime, the algebraic decay yields a derivative scaling of $\frac{\partial\mathcal{M}_{i,i+r}^{(2)}}{\partial V}\sim-2\frac{\partial\alpha}{\partial V}r^{-2\alpha(V)}\ln r$.
-Conversely, in the gapped CDW phase, the exponential decay leads to $\frac{\partial\mathcal{M}_{i,i+r}^{(2)}}{\partial V}\sim\frac{r}{\xi^{2}(V)}e^{-r/\xi(V)}\frac{\partial\xi}{\partial V}$.
+In the gapless LL regime, the algebraic decay yields a derivative scaling of $\frac{\partial\mathcal{M}_{i,i+r}^{(2)}}{\partial V}\sim-2\frac{\partial\eta}{\partial V}r^{-2\eta}\ln r$.
+Conversely, in the gapped CDW phase, the exponential decay leads to $\frac{\partial\mathcal{M}_{i,i+r}^{(2)}}{\partial V}\sim\frac{r}{\xi^{2}}e^{-r/\xi}\frac{\partial\xi}{\partial V}$.
 We identify the location of this maximum as the finite-size critical point, $V_{c}(L)$.
@@
 3(b).
+Notably, analogous finite size scaling behavior has been reported for both density of magic and long range magic (a generalization of the mutual two point SRE) in spin systems [42, 79, 80, 45].
 This result further confirms the capability of the mutual two-point SRE to accurately determine critical exponents at two-dimensional quantum critical points.
@@
 In the CDW phase, the energy gap dictates exponential decay, $\tilde{\mathcal{M}}^{(2)}\sim e^{-r/\xi}$.
-The distinct scaling of the susceptibility $\partial_{V}\tilde{\mathcal{M}}^{(2)}$—diverging as $2\partial_{V}\alpha r^{-2\alpha}\ln r$ in the LL phase versus peaking as $\xi^{-2}re^{-r/\xi}\partial_{V}\xi$ in the CDW phase—signals the phase transition.
+The distinct scaling of the susceptibility $\partial_{V}\tilde{\mathcal{M}}^{(2)}$—diverging as $2(\partial_{V}\eta)r^{-2\eta}\ln r$ in the LL phase versus peaking as $\xi^{-2}re^{-r/\xi}\partial_{V}\xi$ in the CDW phase—signals the phase transition.
 Finite-size effects shift the peak of the latter to $V_{c}(L)$, governed by the BKT-like correlation length scaling $\xi\sim\exp(\pi^{2}/2\sqrt{|V-V_{c}|})$, leading to $V_{c}(L)\approx 2+\pi^{4}/4(\ln L-\ln L_{0})^{2}$.